540 calories of heat converts 1cm cube of water at 100 degrees cecius into 1671 cm cube of steam at 100 degrees celcius at a pressure of 1 atm. then work done against the atmospheric pressure is?

Grade:12

1 Answers

Kirat
6 years ago
Amount of heat given = 540 caloriesChange in volume ∆V =1670 c.cPressure = 1.01×10^6 dyne/cm²W = P∆V =1.01×10^6 × 1670 ____________________. = 40 calories 4.2 × 10^7